Mechanics structural
[MOJ] Solve for the unknown member forces of the entire truss. OB 10. 5K 10 to 10 + 10 to 2014 Upload Choose a File
Answer #1

FBD of whole truan, A 10 8 10_Efo, GiB llo ho Ery=0 RAIRB = 7+5=12 ... EMA=0 (R$40) - (7x20) + (-1*30) R8=7,25k & RA = 4.75k.

FBD of G, TEG Tro ITGH TAB= TBG = TEG Coras & TFG = -10,25k. TAD = 6.72k Tension ТАС - 4:1 (с СОт ел е TCO = 4.75K comprenato